OpenStudy (anonymous):

mmmmkkkayyy. Slightly bewildered here. (1/2)^-x=1.6... I know this is easier than I am making it,Could someone be so kind to explain this?

OpenStudy (kinggeorge):

\[\left({1 \over 2}\right)^{-x} = \left(\left({1 \over 2}\right)^{-1}\right)^x=2^x\]So we have\[2^x=1.6\]From this, we can take the log base 2 of 1.6 to find x.\[x=\log_2(1.6)\]

OpenStudy (kinggeorge):

To find this, you can calculate\[{\log 1.6 \over \log 2}\]

OpenStudy (anonymous):

so if i was to plug that into a calculator I would get the final answer to look something like .678??

OpenStudy (kinggeorge):

That looks correct to me.

OpenStudy (anonymous):

Okay Thanks for the help bro!

OpenStudy (kinggeorge):

You're welcome.